# RoboArmGS: High-Quality Robotic Arm Splatting via Bézier Curve Refinement

### v1

@@ C.1 Novel-View Synthesis
We first evaluate the static reconstruction quality on the Novel-View Synthesis task. As reported in Tab. 1, RoboArmGS achieves excellent rendering fidelity on both robotic arms. Specifically, for the Universal Robots UR5e, our method attains a PSNR of 34.92 dB and an SSIM of 0.968. Similarly, for the ABB IRB 120, we achieve a PSNR of 35.02 dB and an SSIM of 0.969. These metrics indicate that our Structured Gaussian Binding (SGB) strategy effectively leverages the underlying kinematic mesh to constrain the 3D Gaussians, allowing for precise geometry reconstruction even from monocular input. Qualitative results shown in Fig. 1 further corroborate these findings. Our method renders sharp boundaries and preserves fine-grained texture details, effectively capturing the specific visual characteristics of each robot arm without exhibiting the cloudy artifacts often seen in unconstrained Gaussian splatting.
@@ C.2 Novel-Pose Synthesis
The Novel-Pose Synthesis task evaluates the model’s ability to generate photorealistic images under unseen joint configurations, which is critical for building actionable digital assets. Remarkably, as shown in Tab. 1, RoboArmGS achieves even higher quantitative metrics on this dynamic task, with PSNRs exceeding 36 dB and 37 dB for the UR5e and ABB arms, respectively. The LPIPS scores remain exceptionally low (<0.02<0.02), reflecting high perceptual realism. This superior performance is attributed to our Bézier-based Motion Refiner (BMR), which successfully decouples motion artifacts from appearance learning. By compensating for the residual errors between the idealized URDF and the real-world mechanical motion, BMR ensures that the Gaussians are correctly positioned for any given input pose. Fig. 2 visualizes these results, demonstrating that the rendered arms align perfectly with the target poses while maintaining consistent lighting and occlusion handling, proving the effectiveness of our method for high-fidelity dynamic simulation.

### v2

@@ C.1 Novel-View Synthesis
We first evaluate the static reconstruction quality on the Novel-View Synthesis task. As reported in Tab. 4 (left), RoboArmGS achieves leading rendering fidelity on both robotic arms. Specifically, for the UR5e, our method attains a PSNR of 35.220 dB, surpassing the best baseline (SuGaR) by a substantial margin of 4.768 dB. Similarly, for the ABB IRB 120, we achieve a PSNR of 34.815 dB, outperforming Robo-GS by 4.628 dB. These metrics indicate that our Structured Gaussian Binding (SGB) strategy effectively leverages the underlying kinematic mesh to provide strong geometric constraints for the 3D Gaussians. Unlike unconstrained methods like 2DGS or SuGaR, SGB allows for precise geometry reconstruction even from sparse monocular inputs. Qualitative results (see Fig. 6) further corroborate these findings: our method renders sharp boundaries and preserves fine-grained texture details, effectively capturing the specific material characteristics of each robot without the ”cloudy” floaters or structural artifacts often seen in previous Gaussian-based reconstructions.
@@ C.2 Novel-Pose Synthesis
The Novel-Pose Synthesis task evaluates the model’s ability to generate photorealistic images under unseen joint configurations, which is critical for high-fidelity digital assets simulation. Remarkably, as shown in Tab. 4 (right), RoboArmGS achieves even more better results on this dynamic task. For the UR5e arm, our method achieves a PSNR of 39.559 dB, which is a massive 8.6 dB improvement over the second-best method (Deformable 3DGS, 30.959 dB). On the ABB arm, we also maintain a significant lead with 36.379 dB PSNR. The LPIPS scores for both scenes remain exceptionally low, reflecting near-perfect perceptual realism. This superior performance is primarily attributed to our Bézier-based Motion Refiner (BMR). While methods like 4DGS and Deformable 3DGS struggle with large-range articulated movements and cumulative joint errors, BMR successfully compensates for the residual offsets between the idealized URDF model and real-world mechanical motion. By decoupling these motion artifacts from appearance learning, BMR ensures that Gaussians remain accurately attached to their respective links during complex movements. As visualized in Fig. 7, our rendered arms align perfectly with the target poses while maintaining consistent lighting and occlusion handling, proving that RoboArmGS is a highly robust solution for dynamic robotic simulation.